OpenStudy (anonymous):

1. Write an equation in point-slope form for the line through the given points with the given slope. (8,3) m=6 a. y+3=6(x-8) b. y-3=6(x-8) c. y-3=6(x+8) d. y+3=6x+8 2. Graph the equation- y+5=-2(x-4)

OpenStudy (anonymous):

The equation for point-slope form is y-y1=m(x-x1), if you plug in the information provided, you'd get: b. y-3=6(x-8)

OpenStudy (anonymous):



OpenStudy (anonymous):

@amlim14 thanks! can you help with a few more?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

3. which point is located on the line represented by the equation y+4=-5(x-3) a. (-4,-5) b. (-5,-4) c. (3,-4) d. (-3,4)

OpenStudy (anonymous):

@amlim14

OpenStudy (anonymous):

Same thing, just follow the format: y-y1=m(x-x1) y+4=-5(x-3) y-(-4)=-5(x-3), so x=3, y=-4, m=-5. You just need to return it to the original format.